OpenStudy (anonymous):

A aluminium bar has a mass of 60 grams and a molar heat capacity , Cm,p of 24.3 JK-1mol-1 is heated to a temperature of 75 degrees Celsius, and dropped into a bath of water at 15 degrees Celsius. The final temperature of the aluminium bar and bath of water is 20 degrees Celsius. Given that the molar heat capacity of water is 75.3 JK-1mol-1, How many moles of water are contained in the bath? I know the equation is Delta H = mCm,p * Delta T

OpenStudy (aaronq):

find the mass of water in the bath then convert that to moles

OpenStudy (anonymous):

But where do i start? Can you rearrange the equation accordingly?

OpenStudy (aaronq):

You have to assume that all the heat that is lost by the aluminum bar is the heat gained by the water. And, since you're using MOLAR heat capacity, you have to use moles (you need to address (i.e. cancel) the units JK-1mol-1). so, the equation is dH=n*(Cp,m)*dT and since -dH(aluminium) = dH(water), - [n*(Cp,m)*dT] = n*(Cp,m)*dT

OpenStudy (anonymous):

Could you write out your working out please?

OpenStudy (anonymous):

Im still not sure on it :/

OpenStudy (aaronq):

-[(60g/26.98g/mol)*(24.3 J/K*mol)*(20-75)]=(m/18.02g/mol)*(75.3 J/K*mol)*(20-15)

OpenStudy (anonymous):

Were still trying to find m? As on the left I get 2972.2, and at the other side of the = i get 6784.53. Do i subtract these amounts, or use the first as m in the second?

OpenStudy (aaronq):

no sorry i wrote in in mass/molar mass, you could've kept it as n, simply. the answer is 7.89 moles